# StimuVAR: Spatiotemporal Stimuli-aware Video Affective Reasoning with Multimodal Large Language Models

### v1

@@ Abstract
Predicting and reasoning how a video would make a human feel is crucial for developing socially intelligent systems. Although Multimodal Large Language Models (MLLMs) have shown impressive video understanding capabilities, they tend to focus more on the semantic content of videos, often overlooking emotional stimuli. Hence, most existing MLLMs fall short in estimating viewers’ emotional reactions and providing plausible explanations. To address this issue, we propose StimuVAR, a spatiotemporal Stimuli-aware framework for Video Affective Reasoning (VAR) with MLLMs. StimuVAR incorporates a two-level stimuli-aware mechanism: frame-level awareness and token-level awareness. Frame-level awareness involves sampling video frames with events that are most likely to evoke viewers’ emotions. Token-level awareness performs tube selection in the token space to make the MLLM concentrate on emotion-triggered spatiotemporal regions. Furthermore, we create VAR instruction data to perform affective training, steering MLLMs’ reasoning strengths towards emotional focus and thereby enhancing their affective reasoning ability. To thoroughly assess the effectiveness of VAR, we provide a comprehensive evaluation protocol with extensive metrics. StimuVAR is the first MLLM-based method for viewer-centered VAR. Experiments demonstrate its superiority in understanding viewers’ emotional responses to videos and providing coherent and insightful explanations.
@@ 1 Introduction
Understanding human emotional responses to videos is crucial for developing socially intelligent systems that enhance human-computer interaction [1, 2], personalized services [3, 4], and more. In recent years, user-generated videos on social media platforms have become an integral part of modern society. With increasing concerns about mental health, there is growing public attention on how videos affect viewers’ well-being. Unlike most existing Video Emotion Analysis (VEA) approaches that focus on analyzing the emotions of characters in a video [5, 6, 7, 8, 9, 10], predicting and reasoning about a video’s emotional impact on viewers is a more challenging task [11, 12, 13, 14]. This challenge requires not only an understanding of video content but also commonsense knowledge of human reactions and emotions.
@@
Traditional emotion models are trained to map visual embeddings to corresponding emotion labels [5, 6, 7, 8, 13, 15, 16]. These models heavily rely on basic visual attributes, such as color, brightness, or object class [17, 18], which are often insufficient for accurately estimating viewers’ emotional reactions. While recent advances in Multimodal Large Language Models (MLLMs) [19, 20, 21, 22, 23, 24, 25] have demonstrated superiority in various video understanding tasks [26, 27, 28], they tend to focus more on the semantic content and factual analysis of videos. This lack of awareness of emotional knowledge often leads these MLLMs to fall short in viewer-centered VEA.
@@
Psychologists have highlighted that emotions are often triggered by specific elements, referred to as emotional stimuli [29, 30, 31, 32]. Fig. 1 illustrates an example. Consider a 20-second dashcam video in which a rock suddenly falls onto the road within 2 seconds, while the other 18 seconds depict regular driving scenes. Although most of the video depicts ordinary scenes, the unexpected rock fall is likely to evoke surprise and fear in viewers. In this case, the falling rock is a critical stimulus that predominantly shapes the viewers’ emotional responses. However, current MLLMs may overlook or not prioritize these key emotional stimuli. From a temporal perspective, most MLLMs use uniform temporal downsampling to sample input video frames. While uniform sampling may work well for general video understanding tasks, it could miss the unexpected moments or rapid events that could generate strong reactions and thus make a wrong prediction. From a spatial perspective, emotional stimuli like the falling rock may occupy only a small region of the frame. Identifying these stimulus regions is essential for reducing redundant information and achieving more precise affective understanding.
@@
On the other hand, interpretability is crucial for earning public trust when deploying models in real-world applications. Still, traditional emotion models are not explainable, and most current MLLMs fail to provide plausible affective explanations due to their limited awareness of emotional stimuli, as previously discussed. Although a few recent efforts aim at explainable emotion analysis [9, 10, 11, 12, 17], they consider only image data or lack a comprehensive evaluation protocol to fully validate their reasoning ability. The task of reasoning human affective responses triggered by videos remains less explored.
@@
To address these limitations, we propose StimuVAR, a spatiotemporal Stimuli-aware framework for Video Affective Reasoning (VAR) with MLLMs. StimuVAR incorporates a two-level stimuli-aware mechanism to identify spatiotemporal stimuli: frame-level awareness and token-level awareness. For frame-level awareness, we introduce event-driven frame sampling, using optical flow as a cue to capture the frames that contain unexpected events or unintentional accidents [33]. These frames are likely to be the stimuli that evoke viewers’ emotions. For token-level awareness, we design emotion-triggered tube selection that localizes the emotion-triggered spatiotemporal regions in the token space, which the MLLM can then emphasize. In addition, we create VAR visual instruction data via GPT [34, 35] to perform affective training. The VAR-specific instruction data steer the MLLM’s reasoning strengths and commonsense knowledge towards an emotional focus, enhancing the MLLM’s ability to provide insightful and contextually relevant explanations for its affective understanding. To the best of our knowledge, the proposed StimuVAR is the first MLLM-based method for predicting and reasoning viewers’ emotional reactions to videos.
@@
To thoroughly assess the VAR problem, we provide a comprehensive evaluation protocol with extensive metrics, including prediction accuracy, emotional-alignment [11], doubly-right [36], CLIPScore [37], and LLM-as-a-judge [38]. The proposed StimuVAR achieves state-of-the-art performance across multiple benchmarks, demonstrating its ability to predict how a video would make a human feel and to provide plausible explanations. In summary, the main contributions of this work are as follows:
@@ 2.1 Video Emotion Analysis
Existing VEA studies can be categorized into two main themes: recognizing the emotions of characters in a video and predicting viewers’ emotional responses to a video. The former primarily relies on facial expressions and dialogue information [5, 6, 7, 8]. We target the latter, which is more challenging as it requires not only video content understanding but also commonsense knowledge of human reactions. Such viewer-centered VEA is crucial for understanding how a video affects human mental health, gaining more attention in the social media era. Traditional supervised approaches learn to map visual features to corresponding emotion labels [13, 15, 16]. They heavily rely on basic visual attributes, such as color, brightness, or object class [17], which are often inadequate for the complex task of viewer-centered VEA. Additionally, these approaches lack interpretability, failing to provide rationales behind their predictions for broader applications. ArtEmis [11] and Affection [12] introduce benchmarks for viewer-centered affective explanations but only consider image data. Several recent studies explore utilizing MLLMs in emotion analysis. EmoVIT [17] performs emotion visual instruction tuning [39, 40] to learn emotion-specific knowledge. Although EmoVIT claims to have affective reasoning ability, it does not provide an evaluation. EMER [9] and Emotion-LLaMA [10] use MLLMs to offer explanations for their emotion predictions, but they focus on analyzing characters’ emotions rather than viewers’ emotional responses.
@@ 2.2 Multimodal Large Language Models
Recent advances in Large Language Models (LLMs), e.g., the GPT [34, 35] and LLaMA families [41, 42, 43], have demonstrated outstanding capabilities in understanding and generating natural languages. These models have been extended to MLLMs [44, 45, 39, 40] to tackle vision-language problems. For instance, LLaVA [39] and InstructBLIP [40] use a projector to integrate visual features into the LLM and conduct visual instruction tuning for instruction-aware vision-language understanding. Several video-oriented MLLMs have also been introduced specifically for video tasks, such as Video-LLaMA [19], Video-ChatGPT [22], Chat-UniVi [24], mPLUG-Owl [25], etc. However, these approaches typically employ uniform temporal downsampling to sample input video frames, which often results in missing key stimulus moments that are most likely to trigger viewers’ emotions. In contrast, the proposed StimuVAR incorporates spatiotemporal stimuli awareness, enabling it to achieve state-of-the-art VAR performance.
@@ 3.1.1 Event-driven Frame Sampling
In video tasks, uniformly sampling frames is a common practice to represent a video due to temporal redundancy. However, uniform sampling often fails to represent videos containing rapid, unexpected actions or unintentional accidents [33], most likely to evoke viewers’ emotional reactions [32]. This is because uniform sampling could easily miss the frames of such rapid but key events. While processing all frames without sampling would preserve all temporal information, the computational burden is significant, especially for MLLMs. To achieve frame-level stimuli awareness, we aim to develop a sampling method that selects the most representative frames within the same constrained number as the uniform sampling baseline. For practical use, we suggest that such frame sampling should meet the following criteria:
@@
Epstein et al. [33] study a related task of localizing unintentional actions in videos, exploring cues such as speed and context. Still, their approach requires additional deep networks to extract these video features, making it less ideal for incorporation into the MLLM pipeline when considering processing speed.
@@
We propose our event-driven frame sampling strategy based on the observation that rapid key events often coincide with dramatic changes in a video’s appearance. These appearance changes can be modeled using the optical flow estimation technique, widely used to capture motions in video tasks [46, 47]. Let us consider a video $V$ consisting of frames $\{f_{1},f_{2},...f_{T}\}$; an optical flow estimator derives the pattern of apparent motion between each pair of adjacent frames $f_{t}$ and $f_{t+1}$ as follows:
@@
where $OF_{t}$ is a frame-level optical flow value (i.e., mean absolute of each pixel’s optical flow value), and then we obtain a set of estimated optical flows $\{OF_{1},OF_{2},\ldots,OF_{T-1}\}$ of the video. Although there have been many deep network-based optical flow estimators in the literature [48, 49], we find that the classic Lucas-Kanade method [50] is sufficient for our purpose and offers much higher computational efficiency. As illustrated in Fig. 3, we construct a curve $OF(t)$ that depicts the intensity of the optical flows over time. To mitigate noise-induced fluctuations, we apply Gaussian smoothing to this curve using a Gaussian filter $G_{\sigma}$:
@@
Given the higher likelihood of important information within the event-related frames, we assign a high-intensity sampling rate to them, while the non-event frames are sampled with a low-intensity rate. Let us consider a predefined number of frames to sample as $N$, these $N$ frames are evenly distributed across all “events” $\textbf{e}=\{e_{1},e_{2},\ldots,e_{p+1}\}$, where we regard non-event as a single “event”. That is, we uniformly sample $\frac{N}{p+1}$ frames from each event set, and more frames are involved in the non-event set compared to the event set, thereby achieving varying sampling rates. This enables discriminative sampling based on event occurrence, ensuring an efficient allocation of resources to effectively capture the essential frames from the video.
@@ 3.1.2 Emotion-triggered Tube Selection
Inspired by patch selection in visual recognition [51], we formulate token selection as a Top-K problem. As illustrated in Fig. 2, to identify emotional stimuli areas in the token space, we focus on the <Patch> tokens of the visual tokens after the projector, denoted as $\textbf{k}\in\mathbb{R}^{N\times L\times C}$, where $C$ is the embedding dimension and $L$ is the number of tokens. We use a two-layer perceptron $\mathcal{S}$ to estimate correlation scores between each token and the output response, given by $\textbf{c}=\mathcal{S}(\textbf{k})\in\mathbb{R}^{N\times L}$. Next, we reshape the correlation scores c into a 3D volume $\textbf{c}^{r}\in\mathbb{R}^{N\times\sqrt{L}\times\sqrt{L}}$ according to the patch coordinates of the original input frames. This volume is split into tubes of shape $t\times h\times w$, resulting in $((\frac{N-t}{d_{t}}+1)\times(\frac{\sqrt{L}-h}{d_{h}}+1)\times(\frac{\sqrt{L}-w}{d_{w}}+1))$ tubes, where $(d_{t},d_{w},d_{h})$ denotes the stride shape. The token scores within each tube are averaged to obtain each tube’s score, and then we select the Top-K tubes with the highest correlation scores as the final video representation sent to the following LLM.
This design differs from existing token selection approaches [52] that consider the temporal and spatial domains sequentially. In these approaches, temporal selection highly influences the final performance, especially after the frame sampling step. Missing essential frames that trigger emotions during the token selection process significantly reduces the chance of making a correct prediction. We provide empirical evidence in Fig. 6 to support this claim. In contrast, our tube selection strategy groups tokens into tubes, ensuring that the selection is driven by both temporal and spatial information. This design preserves each frame’s intrinsic spatial structure and accounts for the entire video’s consistency and continuity. Furthermore, our tube selection allows for efficient utilization of computational resources while identifying key emotional stimuli in a video.

### v2

@@ Abstract
Predicting and reasoning how a video would make a human feel is crucial for developing socially intelligent systems. Although Multimodal Large Language Models (MLLMs) have shown impressive video understanding capabilities, they tend to focus more on the semantic content of videos, often overlooking emotional stimuli. Hence, most existing MLLMs fall short in estimating viewers’ emotional reactions and providing plausible explanations. To address this issue, we propose StimuVAR, a spatiotemporal Stimuli-aware framework for Video Affective Reasoning (VAR) with MLLMs. StimuVAR incorporates a two-level stimuli-aware mechanism: frame-level awareness and token-level awareness. Frame-level awareness involves sampling video frames with events that are most likely to evoke viewers’ emotions. Token-level awareness performs tube selection in the token space to make the MLLM concentrate on emotion-triggered spatiotemporal regions. Furthermore, we create VAR instruction data to perform affective training, steering MLLMs’ reasoning strengths towards emotional focus and thereby enhancing their affective reasoning ability. To thoroughly assess the effectiveness of VAR, we provide a comprehensive evaluation protocol with extensive metrics. StimuVAR is the first MLLM-based method for viewer-centered VAR. Experiments demonstrate its superiority in understanding viewers’ emotional responses to videos and providing coherent and insightful explanations. Our code is available at https://github.com/EthanG97/StimuVAR.
@@ 1 Introduction
Understanding human emotional responses to videos is crucial for developing socially intelligent systems that enhance human-computer interaction [pantic2005affective, wang2020emotion], personalized services [bielozorov2019role, lee2017chatbot], and more. In recent years, user-generated videos on social media platforms have become an integral part of modern society. With increasing concerns about mental health, there is growing public attention on how videos affect viewers’ well-being. Unlike most existing Video Emotion Analysis (VEA) approaches that focus on analyzing the emotions of characters in a video [srivastava2023you, kosti2019context, lee2019context, yang2024robust, lian2023explainable, cheng2024emotion], predicting and reasoning about a video’s emotional impact on viewers is a more challenging task [achlioptas2021artemis, achlioptas2023affection, mazeika2022would, jiang2014predicting]. This challenge requires not only an understanding of video content but also commonsense knowledge of human reactions and emotions.
@@
Traditional emotion models are trained to map visual embeddings to corresponding emotion labels [srivastava2023you, kosti2019context, lee2019context, yang2024robust, mazeika2022would, yang2021stimuli, pu2023going]. These models heavily rely on basic visual attributes, such as color, brightness, or object class [xie2024emovit, yang2023emoset], which are often insufficient for accurately estimating viewers’ emotional reactions. While recent advances in Multimodal Large Language Models (MLLMs) [Zhang2023VideoLLaMAAI, Lin2023VideoLLaVALU, luo2023valley, Maaz2023VideoChatGPTTD, li2024mvbench, Jin_2024_CVPR, Ye_2024_CVPR] have demonstrated superiority in various video understanding tasks [mittal2024can, Wang2023VamosVA, Yang2024FollowTR], they tend to focus more on the semantic content and factual analysis of videos. This lack of awareness of emotional knowledge often leads these MLLMs to fall short in viewer-centered VEA.
@@
Psychologists have highlighted that emotions are often triggered by specific elements, referred to as emotional stimuli [mehrabian1974approach, frijda1986emotions, brosch2010perception, peng2016emotions]. Fig. 1 illustrates an example. Consider a 20-second dashcam video in which a rock suddenly falls onto the road within 2 seconds, while the other 18 seconds depict regular driving scenes. Although most of the video depicts ordinary scenes, the unexpected rock fall is likely to evoke surprise and fear in viewers. In this case, the falling rock is a critical stimulus that predominantly shapes the viewers’ emotional responses. However, current MLLMs may overlook or not prioritize these key emotional stimuli. From a temporal perspective, most MLLMs use uniform temporal downsampling to sample input video frames. While uniform sampling may work well for general video understanding tasks, it could miss the unexpected moments or rapid events that could generate strong reactions and thus make a wrong prediction. From a spatial perspective, emotional stimuli like the falling rock may occupy only a small region of the frame. Identifying these stimulus regions is essential for reducing redundant information and achieving more precise affective understanding.
@@
On the other hand, interpretability is crucial for earning public trust when deploying models in real-world applications. Still, traditional emotion models are not explainable, and most current MLLMs fail to provide plausible affective explanations due to their limited awareness of emotional stimuli, as previously discussed. Although a few recent efforts aim at explainable emotion analysis [lian2023explainable, cheng2024emotion, achlioptas2021artemis, achlioptas2023affection, xie2024emovit], they consider only image data or lack a comprehensive evaluation protocol to fully validate their reasoning ability. The task of reasoning human affective responses triggered by videos remains less explored.
@@
To address these limitations, we propose StimuVAR, a spatiotemporal Stimuli-aware framework for Video Affective Reasoning (VAR) with MLLMs. StimuVAR incorporates a two-level stimuli-aware mechanism to identify spatiotemporal stimuli: frame-level awareness and token-level awareness. For frame-level awareness, we introduce event-driven frame sampling, using optical flow as a cue to capture the frames that contain unexpected events or unintentional accidents [epstein2020oops]. These frames are likely to be the stimuli that evoke viewers’ emotions. For token-level awareness, we design emotion-triggered tube selection that localizes the emotion-triggered spatiotemporal regions in the token space, which the MLLM can then emphasize. In addition, we construct VAR visual instruction data (based on the training set of the large-scale viewer-centered Video Cognitive Empathy (VCE) dataset [mazeika2022would]) via GPT [brown2020language, achiam2023gpt] to perform affective training. The VAR-specific instruction data steer the MLLM’s reasoning strengths and commonsense knowledge towards an emotional focus, enhancing the MLLM’s ability to provide insightful and contextually relevant explanations for its affective understanding. To the best of our knowledge, the proposed StimuVAR is the first MLLM-based method for predicting and reasoning viewers’ emotional reactions to videos.
@@
To thoroughly assess the VAR problem, we provide a comprehensive evaluation protocol with extensive metrics, including prediction accuracy, emotional-alignment [achlioptas2021artemis], doubly-right [mao2023doubly], CLIPScore [hessel2021clipscore], and LLM-as-a-judge [zheng2024judging]. The proposed StimuVAR achieves state-of-the-art performance across multiple benchmarks, demonstrating its ability to predict how a video would make a human feel and to provide plausible explanations. In summary, the main contributions of this work are as follows:
@@ 2.1 Video Emotion Analysis
Existing VEA studies can be categorized into two main themes: recognizing the emotions of characters in a video and predicting viewers’ emotional responses to a video. The former primarily relies on facial expressions and dialogue information [srivastava2023you, kosti2019context, lee2019context, yang2024robust]. We target the latter, which is more challenging as it requires not only video content understanding but also commonsense knowledge of human reactions. Such viewer-centered VEA is crucial for understanding how a video affects human mental health, gaining more attention in the social media era. Traditional supervised approaches learn to map visual features to corresponding emotion labels [mazeika2022would, yang2021stimuli, pu2023going]. They heavily rely on basic visual attributes, such as color, brightness, or object class [xie2024emovit], which are often inadequate for the complex task of viewer-centered VEA. Additionally, these approaches lack interpretability, failing to provide rationales behind their predictions for broader applications. ArtEmis [achlioptas2021artemis] and Affection [achlioptas2023affection] introduce benchmarks for viewer-centered affective explanations but only consider image data. Several recent studies explore utilizing MLLMs in emotion analysis. EmoVIT [xie2024emovit] performs emotion visual instruction tuning [liu2024visual, Dai2023InstructBLIPTG] to learn emotion-specific knowledge. Although EmoVIT claims to have affective reasoning ability, it does not provide an evaluation. EMER [lian2023explainable] and Emotion-LLaMA [cheng2024emotion] use MLLMs to offer explanations for their emotion predictions, but they focus on analyzing characters’ emotions rather than viewers’ emotional responses.
@@ 2.2 Multimodal Large Language Models
Recent advances in Large Language Models (LLMs), e.g., the GPT [brown2020language, achiam2023gpt] and LLaMA families [touvron2023llama1, touvron2023llama, dubey2024llama], have demonstrated outstanding capabilities in understanding and generating natural languages. These models have been extended to MLLMs [li2023blip, zhu2024minigpt, liu2024visual, Dai2023InstructBLIPTG] to tackle vision-language problems. For instance, LLaVA [liu2024visual] and InstructBLIP [Dai2023InstructBLIPTG] use a projector to integrate visual features into the LLM and conduct visual instruction tuning for instruction-aware vision-language understanding. Several video-oriented MLLMs have also been introduced specifically for video tasks, such as Video-LLaMA [Zhang2023VideoLLaMAAI], Video-ChatGPT [Maaz2023VideoChatGPTTD], Chat-UniVi [Jin_2024_CVPR], mPLUG-Owl [Ye_2024_CVPR], etc. However, these approaches typically employ uniform temporal downsampling to sample input video frames, which often results in missing key stimulus moments that are most likely to trigger viewers’ emotions. In contrast, the proposed StimuVAR incorporates spatiotemporal stimuli awareness, enabling it to achieve state-of-the-art VAR performance.
@@ 3.1.1 Event-driven Frame Sampling
In video tasks, uniformly sampling frames is a common practice to represent a video due to temporal redundancy. However, uniform sampling often fails to represent videos containing rapid, unexpected actions or unintentional accidents [epstein2020oops], most likely to evoke viewers’ emotional reactions [peng2016emotions]. This is because uniform sampling could easily miss the frames of such rapid but key events. While processing all frames without sampling would preserve all temporal information, the computational burden is significant, especially for MLLMs. To achieve frame-level stimuli awareness, we aim to develop a sampling method that selects the most representative frames within the same constrained number as the uniform sampling baseline. For practical use, we suggest that such frame sampling should meet the following criteria:
@@
Epstein et al. [epstein2020oops] study a related task of localizing unintentional actions in videos, exploring cues such as speed and context. Still, their approach requires additional deep networks to extract these video features, making it less ideal for incorporation into the MLLM pipeline when considering processing speed.
@@
We propose our event-driven frame sampling strategy based on the observation that rapid key events often coincide with dramatic changes in a video’s appearance. These appearance changes can be modeled using the optical flow estimation technique, widely used to capture motions in video tasks [turaga2009unsupervised, turaga2007videos]. Let us consider a video $V$ consisting of frames $\{f_{1},f_{2},...f_{T}\}$; an optical flow estimator derives the pattern of apparent motion between each pair of adjacent frames $f_{t}$ and $f_{t+1}$ as follows:
@@
where $OF_{t}$ is a frame-level optical flow value (i.e., mean absolute of each pixel’s optical flow value), and then we obtain a set of estimated optical flows $\{OF_{1},OF_{2},\ldots,OF_{T-1}\}$ of the video. Although there have been many deep network-based optical flow estimators in the literature [dosovitskiy2015flownet, ilg2017flownet], we find that the classic Lucas-Kanade method [lucas1981iterative] is sufficient for our purpose and offers much higher computational efficiency. As illustrated in Fig. 3, we construct a curve $OF(t)$ that depicts the intensity of the optical flows over time. To mitigate noise-induced fluctuations, we apply Gaussian smoothing to this curve using a Gaussian filter $G_{\sigma}$:
@@
Given the higher likelihood of important information within the event-related frames, we assign a high-intensity sampling rate to them, while the non-event frames are sampled with a low-intensity rate. Let us consider a predefined number of frames to sample as $N$, these $N$ frames are evenly distributed across all “events” $\textbf{e}=\{e_{1},e_{2},\ldots,e_{p+1}\}$, where we regard non-event as a single “event”. That is, we uniformly sample $\frac{N}{p+1}$ frames from each event set, and more sampled frames are involved in the event set compared to the non-event set, thereby achieving varying sampling rates. This enables discriminative sampling based on event occurrence, ensuring an efficient allocation of resources to effectively capture the essential frames from the video.
@@ 3.1.2 Emotion-triggered Tube Selection
Inspired by patch selection in visual recognition [cordonnier2021differentiable], we formulate token selection as a Top-K problem. As illustrated in Fig. 2, to identify emotional stimuli areas in the token space, we focus on the <Patch> tokens of the visual tokens after the projector, denoted as $\textbf{k}\in\mathbb{R}^{N\times L\times C}$, where $C$ is the embedding dimension and $L$ is the number of tokens for each frame. We use a two-layer perceptron $\mathcal{S}$ to estimate correlation scores between each token and the output response, given by $\textbf{c}=\mathcal{S}(\textbf{k})\in\mathbb{R}^{N\times L}$. Next, we reshape the correlation scores c into a 3D volume $\textbf{c}^{r}\in\mathbb{R}^{N\times\sqrt{L}\times\sqrt{L}}$ according to the patch coordinates of the original input frames. This volume is split into tubes of shape $t\times h\times w$, resulting in $((\frac{N-t}{d_{t}}+1)\times(\frac{\sqrt{L}-h}{d_{h}}+1)\times(\frac{\sqrt{L}-w}{d_{w}}+1))$ tubes, where $(d_{t},d_{w},d_{h})$ denotes the stride shape. The token scores within each tube are averaged to obtain each tube’s score, and then we select the Top-K tubes with the highest correlation scores as the final video representation sent to the following LLM.
This design differs from existing token selection approaches [wang2022efficient] that consider the temporal and spatial domains sequentially. In these approaches, temporal selection highly influences the final performance, especially after the frame sampling step. Missing essential frames that trigger emotions during the token selection process significantly reduces the chance of making a correct prediction. We provide empirical evidence in Fig. 6 to support this claim. In contrast, our tube selection strategy groups tokens into tubes, ensuring that the selection is driven by both temporal and spatial information. This design preserves each frame’s intrinsic spatial structure and accounts for the entire video’s consistency and continuity. Furthermore, our tube selection allows for efficient utilization of computational resources while identifying key emotional stimuli in a video.